High-speed stroboscopic photographs show that the head of a 260-g golf club is traveling at 46 m/s just before it strikes a 46-g golf ball at rest on a tee. After the collision, the club head travels (in the same direction) at 35 m/s. Find the speed of the golf ball just after impact.
Answer #1

given data :

M = 260

V1 = 46

V2 = 35

m = 46

V = M(V1 - V2)/m = 260(46 - 35)/46 = 62.17 m/s